3x squared,...evaluate this expression for 3 diffrent values of x <------How can I solve that?
Answer by THANApHD(104)

actually I don't know whether U'r quation is (3x)^2 or 3*(x)^2
3 x squared can be written like 3*x^2
So evaluvate x's 3 different values mean we have to substitute three different numerical values for x in the equation.
x=(1,0.98,1/5,100 ... ) as a variable

for an eg: 3*(2)^2= 3*4 = 12
3*(1/2)^2 = 3 * 1/4 = 3/4
3 * (10)^2 = 3 * 100 = 300